### v1

# Balancing Confidentiality and Transparency for Blockchain-based Process-Aware Information Systems

Abstract
Blockchain enables novel, trustworthy Process-Aware Information Systems (PAISs) by enforcing the security, robustness, and traceability of operations. In particular, transparency ensures that all information exchanges are openly accessible, fostering trust within the system. Although this is a desirable property to enable notarization and auditing activities, it also represents a limitation for such cases where confidentiality is a requirement since interactions involve sensible data. Current solutions rely on obfuscation techniques or private infrastructures, hindering the enforcing capabilities of smart contracts and the public verifiability of transactions. Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs aimed at preserving both confidentiality and transparency. Smart contracts enact, enforce and store public interactions, while attribute-based encryption techniques are adopted to specify access grants to confidential information. We assess the security of our solution through a systematic threat model analysis and assess its practical feasibility by gauging the performance of our implemented prototype in different scenarios from the literature.
Keywords:
Business process management Distributed ledger technologies Blockchain Attribute-based encryption Security Privacy1 Introduction
Blockchain enables new forms of trustable Process-Aware Information Systems (PAISs) and inter-organizational collaborations [36]. This is possible thanks to the characteristics of public permissionless blockchains, which provide strong guarantees removing the need for third-party authorities [30, 31]. Starting from the foundation, the distributed ledger and consensus mechanisms permit distributed and immutable information storage. These, combined with the cryptographic primitives, ensure security properties and accountability of recorded operations. Smart contracts are programs immutably stored inside the blockchain that can encode and enforce business logic. On top of this, transparency is a key aspect, as information stored in the ledger and performed operations are accessible to everyone within the system.
Business logic enforcement and transparency are highly desirable properties for collaborative process execution and monitoring [13, 14]. However, public observability of the whole set of exchanged data hinders the adoption of blockchain-based solutions in contexts where confidentiality is a critical requirement. Striking a balance between these two aspects is crucial in highly regulated sectors wherein sensitive data are treated (consider, for instance, the domains of pharmaceutical supply-chain and healthcare [18, 2]). Existing solutions to solve this conundrum employ private permissioned blockchains [39, 7]. However, these solutions may hide data from auditors and assume high robustness and security guarantees of the consortium’s system. Other approaches propose encryption techniques to restrict visibility of data stored on-chain only to authorized parties [28, 29, 23, 24]. However, their integration with process management systems is limited or non-existing.
Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs that preserves confidentiality of exchanged information while keeping public enforcement and transparency of process execution. To this end, we bridge the silos of blockchain technologies and encryption with business process management. In particular, we rely on smart contracts to encode and execute business process logic while logging the interactions between parties, and we resort to Attribute-Based Encryption (ABE) [34] to control the access of different parties to the activities’ data payloads and information artifacts, thus safeguarding confidentiality. We demonstrate the security of our approach against a threat model of the proposed architecture. Following the principles of Design Science Research (DSR), we implemented our artifact and evaluated the feasibility of our solution by showing its application in the context of a healthcare scenario taken from the literature and analyzing the execution performance with collaborative processes presented in the related scientific literature.
The remainder of the paper is organized as follows. Section 2 illustrates the example and highlights the motivation for this work. Section 3 provides the analysis of the state of the art and introduces the fundamental concepts on which CONFETTY relies. Section 4 describes the proposed architecture. Section 5 evaluates our solution and provides an assessment of its security and feasibility. Finally, Sect. 6 concludes the work and draws future research directions.
2 Example and Problem Illustration
To illustrate the problem we aim to tackle, we introduce a running example in the healthcare domain inspired by [9]. The model depicted in Fig. 1 depicts the management process for X-ray diagnostics in the form of a BPMN choreography diagram [11]. The patient first presents the medical prescription, asking for an appointment. The radiology department checks the availability of the ward and, if a date is available, the appointment is confirmed. Otherwise, there is a new scheduling tentative. After obtaining the registration details, the temperature and vaccine certification of the patient are verified. If the temperature is under the threshold and the certification is valid, a radiology clerk checks the appointment identifier of the patient and returns the confirmation. If the appointment is confirmed, the X-ray exam is done and the results are provided to the ward with a final report. Finally, the patient receives the results from the ward.
As the example shows, ensuring accurate tracking of medical procedures is critical to improving process efficiency, trust in the public towards the healthcare system, and enabling auditing in case of misconduct or compliance with regulations. Data-based evidence is crucial to this end. The immutability, accessibility and persistence guarantees offered by a public-blockchain-based PAIS are fit for purpose from this standpoint. However, sensitive patient data must also be kept shielded from unauthorized inspections: publicly disclosing the result of the diagnostic analysis or information about vaccine certification would severely violate the privacy sphere of the patient, for example. Balancing a secure, publicly verifiable and decentralized infrastructure with the need for data confidentiality remains therefore an open challenge, which motivates our research.
3 Background and State of the Art
In this section, we first describe the main concepts on which our approach relies, namely blockchain technology and Multi-Authority Attribute-Based Encryption (MA-ABE), and then illustrate how those pillars constitute the foundations of existing approaches in the literature that are of relevance to our investigation.
3.1 Background
Blockchain technology.
A ledger is an append-only singly-linked list of transactions, namely the transfer of assets among accounts. Every account is linked to a pair of keys: a public one for recognition and addressing, and a public one for the unlocking and expenditure of the account’s balance. A blockchain collates consecutive segments of the ledger in its body, while the header includes, among others, the number of the current block and the hash of the current and the previous block. The hash-link to the previous block preserves the total order of transactions in the ledger [12]. This hash cryptographically ensures that previous blocks, and the data therein, are not changed since this would corrupt the hash-links. Transactions can thus be included in an append-only manner, and it is not possible to delete old interactions. New blocks are added to the chain at regular intervals of time by selected computational nodes within the blockchain infrastructural network. Those nodes are needed to compute and finalize the blocks’ hash codes, among other things. New cryptocurrency is minted and distributed to the nodes’ accounts to pay back their computational effort. The entire blockchain data structure is stored and replicated among all nodes in the network, who have full visibility over the saved ledger. The use of consensus protocols and cryptographic primitives permits the convergence and public verification of the transactions included in every block of each replica, thereby guaranteeing decentralization and trustworthiness without the need for an independent third-party authority certification or governance. Several platforms implement the blockchain protocol (e.g., Ethereum [41]). In the last few years, blockchain technologies have been adopted in many different contexts, especially after the introduction of smart contracts, namely programs deployed onto the blockchain and executed within a dedicated virtual machine (e.g., the Ethereum Virtual Machine, EVM). Their functions are invoked through transactions. Core features of smart contracts are the immutability of their code and the storage of their invocations on the blockchain, making such information available for auditing purposes [15]. A smart contract has an account in which it can store and update data alongside its cryptocurrency balance. Smart contract executions incur a cost on the invoker. The cost is measured in gas units and depends on the complexity of the called function and the size of the exchanged data. To save on the latter, distributed tamper-proof storage systems are often employed to store the actual data in a non-modifiable file, while a permalink to that file is saved in the smart contract state to the datastore’s locator. The InterPlanetary File System (IPFS) [3] is a typical example of such a system. Based on a peer-to-peer network wherein each node stores chunks of the data, IPFS binds every file with a string used both as a unique identifier and resource locator. The string is based on the hash of the file content itself so that if the file is altered, the permalink does not match. The code execution is validated in each node of the blockchain, providing verifiable proof of the obtained results for the final users [20]. As we will see next, these features fostered the development of blockchain-based PAISs using smart contracts as the backbone for business (process) rule enforcing (in our example, the BPMN choreography is translated into smart contract code, and its steps correspond to invocations to its functions) [9, 13].
Attribute-based Encryption.
Attribute-Based Encryption (ABE) is a type of public key encryption where the ciphertext, namely an encrypted plaintext, and the corresponding decryption key are linked through custom boolean attributes [34]. Ciphertext-Policy Attribute-Based Encryption (CP-ABE) [5], a variant of ABE, prescribes that every user is endowed with a set of attributes encoded in their key. For example, user 0xB0…1AA1 can be endowed with attributes Patient to specify their role and attribute PID476948 to indicate they are involved in process instance .
Data owners write policies to be attached to the CP-ABE-encrypted document (the ciphertext). Those policies express conditions for granting in-clear access to the data. Policies are propositional logic formulae employing ABE attributes as literals. The policy is evaluated on the attributes associated with a potential reader: Only if the formula is satisfied, the user’s key can decrypt the data.
For instance, the key of the user 0xB0…1AA1 can decrypt data locked with a policy like PID476948 and PATIENT, but cannot unlock the contents of a document associated with a policy like PID476949 and RADIOLOGY.
Attributes are typically assigned by one authority:
3.2 State of the Art
Over the last few years, several solutions that automate collaborative processes using blockchain technology have been designed [13, 8]. Several works in this field have demonstrated the effectiveness of blockchain-based solutions in improving trust among participants in multi-party collaborations [40] even in adversarial settings [26], enhance verifiability of workflows with model-driven approaches [25, 38], allow for monitoring [15], mining [21], security [23], and auditing [10, 7]. These studies enhance the integration of blockchain technology with process management, unlocking security and traceability benefits. However, they primarily focus on the control-flow perspective and lack mechanisms for secure access control to the stored data on public platforms.
A relevant area of research involves allowing for data privacy and integrity within blockchain systems. Several studies examine the application of encryption techniques to achieve this goal. Hawk [22] employs user-defined private Smart Contracts to automate the implementation of cryptographic protocols. Abid et al. [1] leverage Zero-Knowledge Proof (ZKP) and Fully Homomorphic Encryption (FHE) to build a solution that supports privacy-preserving traceability and the analysis of encrypted data. They store on-chain only the hash of the data used to verify the correctness, while all the computations and the data are stored off-chain. Benhamouda et al. [4] use a public blockchain storage layer for confidential data. Peng et al. [32] propose a consortium blockchain as a decentralized storage system for file sharing among organizations. Wu et al. [42] present an attribute-based access control scheme that preserves the privacy of both attributes and policies. Pham et al. [33] propose a decentralized storage solution that integrates InterPlanetary File System (IPFS), MA-ABE, and blockchain technology. Hong et al. [19] introduce a data-sharing solution based on decentralized ABE (MA-ABE), blockchain, and IPFS. Bramm et al. [6] present BDABE, an access control mechanism that applies a distributed ABE scheme within a consensus-driven infrastructure for real-world applications. Feng et al. [16] combine attribute-based encryption ABE and identity-based encryption (IBE) to achieve efficient data sharing and verification of data correctness. They employ blockchain technology to ensure tamper-proof and regulated data storage. Yan et al. [43] introduce a scheme for fine-grained access control by implementing proxy encryption and decryption while supporting policy hiding and attribute revocation. The encrypted data is stored on IPFS, and the metadata is stored on the blockchain. All these studies, along with the aforementioned [29, 28, 27], allow for secure access control on blockchains but lack the integration of this technology with process management systems.
To the best of our knowledge, this is the first work that focuses on combining blockchain-based process execution engines with encryption schemes that guarantee data confidentiality and access control over data shared on public blockchain platforms. In the next section, we provide a description of our solution.
4 The CONFETTY architecture
In this section, we describe our proposed architecture, named CONFidentiality EnforcemenT TransparencY (CONFETTY). We first introduce its key functionalities and the core components that support them (Sect. 4.1), then describe how these components interact and manage information to support confidential-preserving blockchain-based process execution (Sect. 4.2). Due to space restrictions, we do not enter the technical details of our approach but provide an overarching presentation of our solution’s rationale and key concepts. Finally, we conclude the section with some considerations on the benefits brought by our approach merging blockchain technologies and ABE.
4.1 Functional viewpoint
Figure 2 outlines the CONFETTY architecture,
which caters for the following main functionalities.
(i) The configure functionality foresees the specification of the process by the process owner. It requires the registration of process information including activities to execute, data to exchange, control-flow routing, the expected participants’ roles (e.g., the information enclosed in the choreography diagram in Fig. 1) and how they can operate with confidential data (via ABE policies) to control information access.
(ii) The instantiate functionality lets certifiers authenticate the participants, define their role in the process (e.g., that user 0xB0…1AA1 is a patient), and also kickstart the process run on-chain, by setting its public state.
(iii) The transact functionality allows process participants to interact following the process to exchange data at run time. This functionality manages both public and confidential data, updating and enforcing the process state according to the logic in the blockchain.
(iv) The inspect functionality allows participants and auditors to access public and confidential data.
The functionalities above are realized by four main components of CONFETTY: (i) The Process Manager handles all the operations involving the participants and related to the process. This is done by acting as an interface for the blockchain, abstracting from the technical implementation and exposing functionalities to the users. In particular, the Process Manager instantiates the process on the blockchain on given process specification data. To support the instantiate functionality, it provides capabilities for role assignment to participants. (ii) The Process Contract is a smart contract that manages the process instances and their execution. It elaborates and manages all process specifications and their instances’ public states, acting as a software realization both of factory and proxy patterns [17] akin to [10]. It supports the instantiate functionality by enforcing the control flow, the data flow, the assignment of tasks to participants, and the routing of decision points. In this way, the Process Contract updates the public state of the process and generates transparent records that are accessible and verifiable by interested auditors. (iii) The Confidentiality Manager is in charge of handling the participants’ authorizations and the operations over confidential data. It operates as a facade towards the blockchain to store access grants, distributes decryption keys to the users solely based on their roles and involvement in the process, and handles the storage and retrieval of encrypted data. (iv) The Confidentiality Contract is a smart contract handling the notarization of authorizations for confidential data, as well as for their writing and access. In addition to the above inner components, the architecture relies on two external ones: a Programmable Blockchain Platform (like Ethereum) and a Tamper-proof Distributed File Storage (like IPFS). The former maintains the public state, allowing for the creation of logic and constraints while notarizing the executed operations. It comprises two core elements: The Distributed Ledger, storing the transactions, and the Virtual Machine, hosting and running smart contracts. The latter (which we will henceforth name as Distributed FS for brevity), is used to store business data and thus improve the performance and costs of a blockchain. The data is saved on the external storage (upon encryption, if it must remain secret), while a resource locator with the hash of that data is kept on-chain for notarization and future retrieval.
We conclude this subsection with a few considerations about the actors and trust model. We assume that the aforementioned users (process owners, certifiers, process participants, and auditors) possess a blockchain account with a key pair to be able to publicly identify themselves (as with the aforementioned user 0xB0…1AA1) and sign blockchain transactions built through the CONFETTY modules. We also remark that we consider the CONFETTY platform as a trusted party, while users are assumed to be honest but curious. Finally, notice that the involvement of a pre-appointed third-party certifier to authenticate process participants is necessary as a user alone can confirm their identity (e.g., via the aforementioned public-private key scheme) but cannot attest alone to the truthfulness of assertions regarding themselves.
4.2 Information viewpoint
After providing an overview of the CONFETTY architecture, we describe how information is handled in realizing the aforementioned functionalities by the CONFETTY components.
Figure 3 depicts the configure functionality.
The process owner sends the process specification files (like the diagram in Fig. 1) to the Process Manager (1.0). The Process Manager, in turn, (1.1) processes the input data, storing the behavioral and business constraints it reports. Also, (1.2) it builds a policy template, associating roles of the process with the data they can access via encryption directives. Once created, the policy template is returned to the process owner.
In our running example, an extract of a policy template is:
{medicalPrescription:($PID and (PATIENT or RADIOLOGY)),
registration:
($PID and (RADIOLOGY or PATIENT)), report:($PID and (RADIOLOGY or WARD))}.
Here medicalPrescription, registration, and report represent messages, while PATIENT, RADIOLOGY and WARD are the roles of users who can access them. $PID is a placeholder for the process instance identifier (a piece of information that can be known only at runtime).
We give the process owner the option to revise the templates. In this way, adding custom roles that are not expected active parties in the process but need access to data. This is the case, for instance, of external auditors who need to certify data and its compliance with norms, laws and regulations. Table 1 shows an extract from the final access policies of our running example. In this case, each message is attached to a policy with an ID representing the process instance and the authorized role. Considering the medicalPrescription message, the policy specifies that to access it, a user has to be attested to have the role of a MINISTRY-INSPECTOR, or again to be a PATIENT or RADIOLOGY role within the running process instance.
| Message | Access policy |
| Medical prescription | MINISTRY-INSPECTOR or ($PID and (PATIENT or RADIOLOGY)) |
| Registration | MINISTRY-INSPECTOR or ($PID and (RADIOLOGY or PATIENT)) |
| Report | MINISTRY-INSPECTOR or ($PID and (RADIOLOGY or WARD)) |
Once policies are confirmed, (1.3) they are sent to the Confidentiality Manager, which (1.4) stores them in the Distributed FS. Their permalink (1.5) is passed back to the Confidentiality Manager which saves it.
Figure 4 depicts the instantiate functionality. Differently from the configure functionality, needed once per process, this one is designed for every new instantiation thereof. To initialize a new process instance (e.g., number 476948), participants send their data to the Process Manager (2.0) to authenticate themselves as participants in the process with their role (e.g., that user 0xc5…Fd43 operates as a WARD clerk in a new process instance PID476948). (2.1) The authentication is attested to by certifiers confirming the user’s self-declared roles.
When all process participants are identified and associated with a role, (2.2) the Process Manager merges the received information with the process data acquired in step 1.1. Thereupon, (2.3) a selected participant kickstarts a new process instance through the Process Manager associated with the aforementioned information on the Process Contract, thereby initializing the instance’s public state. Notice that also this crucial step is notarized on chain (2.4). The selection of the participant, or participants, on behalf of whom the transaction is signed, transcends the scope of this paper and depends on the decision process and guidelines of the consortium.
After initializing the process state, (2.5) the Process Manager triggers the Confidentiality Manager, which (2.6) stores the policy locator previously saved on the Confidentiality Contract. Finally, (2.7) the Confidentiality Manager notarizes the operation.
The transact functionality reported in Fig. 5 represents the exchange of information during the execution of an activity in the process instance. This passage updates the public state update in the blockchain according to previous enforcement (e.g., activity to execute, expected participant, etc.) and manages accessibility in case of public and confidential cases.
The first case refers to public data and is displayed in Fig. 5(a).
Initially, (3.0a) a user sends the execution data to the Process Manager in the context of an activity. For example, the WARD clerk identified by 0xc5…Fd43 replies to the
“Check availability” activity request by the RADIOLOGY user within process instance 476948. To do so, the clerk sends a “Check availability” response message wherein the accepted variable is assigned with true since the appointment is confirmed alongside an available date. The boolean answer (accepted) is public data that is used later to make control-flow decisions and advance the state of the process, defining the next activity to perform. In our scenario, the exclusive choice gateway is based on the value of the accepted variable.
Then, (3.1a) the Process Manager invokes the Process Contract to request a state update, sending the execution data.
In turn, (3.2a) the Process Contract enforces the business and control-flow logic through the run of its routine (based on previously agreed upon and immutable code, see step 1.1), which checks if the executed activity is enabled and if the message sender (in this case, user 0xc5…Fd43) is the expected one (a WARD), finally updating the state if the verification outcome is positive.
Also, (3.3a) this step advances the state of the process, updating the expected next activity (here, “Confirm appointment”) and participant (RADIOLOGY). This piece of information is notarized into the Distributed Ledger.
The confidential case is depicted in Fig. 5(b). The Process Manager, (3.0b) once it has received the data, (3.1b) forwards it to the Confidentiality Manager. The latter (3.2b) retrieves the policy from the Distributed FS, and (3.3b) uses it to encrypt the data via Attribute-Based Encryption (ABE). For example, considering Fig. 1 and Table 1, the medicalPrescription of the patient is encrypted with the following policy: {MINISTRY-INSPECTOR or (PID476948 and (PATIENT or RADIOLOGY))}.
In particular, the Confidentiality Manager retrieves the specific authorization for the data to be stored and embeds it in the encryption mechanism, ensuring that only authorized participants can access it.
Once encrypted, (3.4b) the Confidentiality Manager stores data in the Distributed FS, (3.5b) its locator in the Confidentiality Contract, and (3.6b) notarizes the operation in the Distributed Ledger.
In this case, the resource locator that contains the encrypted medicalPrescription (e.g., LXBw…2DAx) is Qmo9…0weI.
The execution of this routine by the Confidentiality Contract includes a call to the Process Contract (3.7b) to request a state update (enclosing the execution data). Following the example, this data is the activity being executed (“Check availability”), the user that performs it (e.g.,, PATIENT, along with their corresponding address 0x2e…6dd9), and the process instance (e.g.,476948).
Then, (3.8b) the Process Contract checks the current public state to update it. The operations involving the smart contracts (3.9b) are notarized in the Distributed Ledger.
The inspect functionality pertains to the access to message data. The public data can be freely accessed, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above). To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester. This is the case, e.g., for the accepted field of the “Check availability” task. Reading confidential data requires more passages, illustrated in Fig. 6, and requires a preliminary step. All users should request at least once (and at every attribute update) a decryption key (Fig. 6(a)), to be used for every message they are going to (Fig. 6). For the first one, namely the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a). The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger. Then, (4.3a) the Confidentiality Manager generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via Attribute-Based Encryption (ABE), and (4.4a) sends it back. The decryption of confidential data is depicted in Fig. 6(b). Once the process participant has received the a-b key, they can use it to decrypt the confidential data. To do so, (4.0b) the user makes a request to the Confidentiality Manager. Then, (4.1b) the Confidentiality Manager retrieves the encrypted data from the Distributed FS and (4.2b) forwards it to the user. At this point, the latter (4.3b) uses the a-b key to decrypt the confidential data via ABE. Only if the user’s attributes satisfy the policy previously used to encrypt the data, can the user’s a-b key decrypt the confidential data (and the user inspect it in-clear).
A note on decryption keys.
We remark that an a-b key stays with the user and lasts for as long as their attributes are not updated (in case, e.g., a user is involved in a new process instance, say 476949, a new attribute gets associated to them: $PID476949). The same key will be used to access all messages. For example, consider user 0xB0…1AA1 (who holds the attributes PID476948 and PATIENT).
As per the policies in Table 1, their a-b key can decrypt both the medical prescription and the registration alike, but not the report.
The a-b key of user 0xc5…Fd43 (PID476948 and PATIENT) can decrypt the report instead, but not the other two documents.
Notice that a HEALTH-INSPECTOR, instead, can use their unique a-b key to access all the aforementioned documents regardless of the process instance they pertain to.
The key, in other words, is associated to the user and it is the sole instrument to access a document.
Therefore, the Confidentiality Manager does not need to generate a different decryption key for every document and give the same copy to multiple users, nor create multiple copies of the same document for different decryption keys (one per user), thus saving on unnecessary data replications and key distributions. Also, as the key stays with the users, and the documents are stored and notarized on tamper-proof external infrastructures (the Distributed FS and Blockchain Platform, respectively) the information remains available even should the CONFETTY system stop functioning. These aspects have an impact on security considerations, as we will discuss next.
5 Evaluation
To assess the feasibility of the CONFETTY architecture we implemented a supporting prototype and evaluated it. We first describe the implemented architecture analyzing its thread model, then we evaluate CONFETTY in operation.
5.1 CONFETTY Implementation and Evaluation
Figure 7 illustrates the technologies utilized in the implementation of CONFETTY. Our prototype is deployed on three main tiers. An application server hosts the Confidentiality Manager and the Process Manager, developed in Python 3.6.9 and Java SDK-13, respectively. Users communicate with it via TLS communication channels [37]. We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache. Accessed: December 7, 2024. as the Distributed Tamper-proof File Storage. We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js.1 The source code of CONFETTY, alongside the experimental data we use, is publicly available at github.com/apwbs/CONFETTY.
5.2 Threat Model
In this section, we identify potential security threats that could compromise data confidentiality within our solution, and outline how the CONFETTY architecture addresses them. Our threat analysis is based on the STRIDE framework [35], a theoretical model that categorizes threats into six distinct groups: spoofing (i.e., impersonation of a legitimate entity), tampering (i.e., unauthorized modification of data to compromise its integrity), repudiation (i.e., the denial of having performed a particular action), information disclosure (i.e., the unauthorized exposure of sensitive data), denial of service (DoS, i.e., the disruption or degradation of system availability), and elevation of privileges (i.e., the unauthorized acquisition of higher-level access rights).
As said in Sect. 4.1, we assume users and components to be honest but curious and trusted, respectively. The possible threats thus reside in the communication between them, subject to potential attacks. Figure 8 displays a Data Flow Diagram (DFD), in which the orange dotted lines indicate the trust boundaries crossing data exchange points between different perimeters, represented by black arrows connecting different elements. The opening side of each boundary indicates the trusted component in the information exchange.
The CONFETTY boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditory). In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented. As for tampering, trust is based on the fairness of the user who sends the data. It cannot be the case for a user to repudiate an action because all of them are recorded on the blockchain with a transaction signed with the user’s private key. An attacker could sniff the data a user sends to CONFETTY. However, for those communications we resort to TLS channels, specifically designed to prevent information leakage [37]. DoS attacks are in principle feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform. Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1.0 in Sect. 4.2) which malicious nodes could exploit with multiple calls and bulky payloads. However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss. Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE. Elevating a user’s privileges is not possible since attributes and permissions are linked to a blockchain account.
The Blockchain Platform is the trusted component in the Blockchain Boundary (BB). In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the real one. This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their own private keys. Information disclosure can be achieved only on public data or resource locators of sensitive data. However, the latter is stored upon ABE-encryption, making its content unreadable by malicious users. Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols.
As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack, and elevation of privileges, are counteracted by design, too. Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain. Since we resort to ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.
5.3 Execution Analysis
[…]
Table 2 shows the execution time and costs registered during the execution of the X-ray diagnostic process. For the time evaluation, we ran a separate process instance for each path in the control flow (each corresponding to a different decision in correspondence with the exclusive-choice gateways) five times (thus totalling twenty executions). The registered values are reported in Table 2(a). We employed the Sepolia testnet as the Blockchain Platform, and passed input data based on real-world files in the healthcare domain upon anonymization. The dataset and the Sepolia transaction handles are openly accessible at github.com/apwbs/CONFETTY.
[…]
| min | max | avg | |
| Configure | |||
| Instantiate | |||
| Transact | |||
| Inspect | |||
| Functionality | Used gas units |
| Instantiate | |
| Transact | |
| Inspect | |
| Deployment | |
| Total |
[…]
The second part of our experimental evaluation illustrates the time and cost needed by CONFETTY to run three different processes from the literature, namely the aforementioned X-ray [9], retail [7] and incident management [10]. Figure 9 illustrates the results, averaging overruns that took different branches in correspondence with exclusive choice gateways. We gathered data using a local RPC as a stub for the Programmable Blockchain Platform. The three processes exhibit similar trends from both standpoints. The Incident management process is the heaviest and most expensive one, yet exhibiting an increment of for the cumulative time and for the cumulative cost with respect to the lower bound of the X-ray process. This increment depends on the number of process participants ( in the Incident management against the of the X-ray process), thus leading to the execution of more steps.
6 Conclusions
In this work, we presented CONFETTY, a blockchain-based architecture for PAISs. CONFETTY allows for transparency and public enforcement of process execution while preserving confidentiality of sensitive data. A public state saved on chain maintains process data used for enforcement and auditability purposes. The evolution of the public state is mediated by smart contract implementing the business process control- and data-flow logic. We use ABE encryption to hide sensitive information and grant access to it only to authorized parties based on their roles in the process, encoded as attributes in access policies and embedded directly in their personal decryption keys. To evaluate the CONFETTY architecture we have analyzed its threat model and measured its time and cost performance through the execution of an X-ray examination choreography process. Future work includes extensive performance and scalability analyses against synthetic and real-world benchmarks. We also envision a field study to assess the practical adoption and usability with real-world stakeholders. Another interesting avenue for further investigations is the full decentralization of the architecture, to remove the existing points of trust residing in off-chain components. Finally, we foresee a study and design of a solution to enable automated decision support via computation held on encrypted data.

### v2

@@
 The inspect functionality pertains to the access to message data.
-The public data can be freely accessed, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above).
+The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above).
 To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester.
@@
 All users should request at least once (and at every attribute update) a decryption key (Fig.
-6(a)), to be used for every message they are going to (Fig.
-6).
-For the first one, namely the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a).
+6(a)), to be used for every message they are going to read (Fig.
+6(b)).
+For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a).
 The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger.
@@
 As per the policies in Table 1, their a-b key can decrypt both the medical prescription and the registration alike, but not the report.
-The a-b key of user 0xc5…Fd43 (PID476948 and PATIENT) can decrypt the report instead, but not the other two documents.
+The a-b key of user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
 Notice that a HEALTH-INSPECTOR, instead, can use their unique a-b key to access all the aforementioned documents regardless of the process instance they pertain to.
@@
 We first describe the implemented architecture analyzing its thread model, then we evaluate CONFETTY in operation.
-5.1 CONFETTY Implementation and Evaluation
+5.1 Implementation
 Figure 7 illustrates the technologies utilized in the implementation of CONFETTY.
@@
 We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache.
-Accessed: December 7, 2024.
+Accessed: December 13, 2024.
 as the Distributed Tamper-proof File Storage.
@@
 The opening side of each boundary indicates the trusted component in the information exchange.
-The CONFETTY boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditory).
+The CONFETTY Boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditor).
 In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented.
@@
 4.2) which malicious nodes could exploit with multiple calls and bulky payloads.
-However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss.
+However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss.
 Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE.

### v3

@@
 In particular, transparency ensures that all information exchanges are openly accessible, fostering trust within the system.
-Although this is a desirable property to enable notarization and auditing activities, it also represents a limitation for such cases where confidentiality is a requirement since interactions involve sensible data.
-Current solutions rely on obfuscation techniques or private infrastructures, hindering the enforcing capabilities of smart contracts and the public verifiability of transactions.
-Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs aimed at preserving both confidentiality and transparency.
+Although this is a desirable property to enable notarization and auditing activities, it also represents a limitation for such cases where confidentiality is a requirement since interactions involve sensitive data.
+Current solutions rely on obfuscation techniques or private infrastructures, hindering the enforcement capabilities of smart contracts and the public verifiability of transactions.
+Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs to preserve confidentiality and transparency.
 Smart contracts enact, enforce and store public interactions, while attribute-based encryption techniques are adopted to specify access grants to confidential information.
-We assess the security of our solution through a systematic threat model analysis and assess its practical feasibility by gauging the performance of our implemented prototype in different scenarios from the literature.
+We assess the security of our solution through a systematic threat model analysis and evaluate its practical feasibility by gauging the performance of our implemented prototype in different scenarios from the literature.
 Keywords:
@@
 Blockchain enables new forms of trustable Process-Aware Information Systems (PAISs) and inter-organizational collaborations [27].
-This is possible thanks to the characteristics of public permissionless blockchains, which provide strong guarantees removing the need for third-party authorities [30, 31].
+This is possible thanks to the characteristics of public permissionless blockchains, which provide strong guarantees removing the need for third-party authorities [22].
 Starting from the foundation, the distributed ledger and consensus mechanisms permit distributed and immutable information storage.
-These, combined with the cryptographic primitives, ensure security properties and accountability of recorded operations.
+These features, combined with the cryptographic primitives, ensure security properties and accountability of recorded operations.
 Smart contracts are programs immutably stored inside the blockchain that can encode and enforce business logic.
 On top of this, transparency is a key aspect, as information stored in the ledger and performed operations are accessible to everyone within the system.
-Business logic enforcement and transparency are highly desirable properties for collaborative process execution and monitoring [13, 14].
+Business logic enforcement and transparency are highly desirable properties for collaborative process execution and monitoring [10].
 However, public observability of the whole set of exchanged data hinders the adoption of blockchain-based solutions in contexts where confidentiality is a critical requirement.
-Striking a balance between these two aspects is crucial in highly regulated sectors wherein sensitive data are treated (consider, for instance, the domains of pharmaceutical supply-chain and healthcare [18, 2]).
+Striking a balance between these two aspects is crucial in highly regulated sectors wherein sensitive data are treated (consider, for instance, the domains of the pharmaceutical supply-chain and healthcare [11]).
 Existing solutions to solve this conundrum employ private permissioned blockchains [30, 6].
-However, these solutions may hide data from auditors and assume high robustness and security guarantees of the consortium’s system.
+Nevertheless, these solutions may hide data from auditors and assume high robustness and security guarantees of the consortium’s system.
 Other approaches propose encryption techniques to restrict visibility of data stored on-chain only to authorized parties [15, 16, 20, 21].
-However, their integration with process management systems is limited or non-existing.
-Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs that preserves confidentiality of exchanged information while keeping public enforcement and transparency of process execution.
-To this end, we bridge the silos of blockchain technologies and encryption with business process management.
-In particular, we rely on smart contracts to encode and execute business process logic while logging the interactions between parties, and we resort to Attribute-Based Encryption (ABE) [34] to control the access of different parties to the activities’ data payloads and information artifacts, thus safeguarding confidentiality.
-We demonstrate the security of our approach against a threat model of the proposed architecture.
-Following the principles of Design Science Research (DSR), we implemented our artifact and evaluated the feasibility of our solution by showing its application in the context of a healthcare scenario taken from the literature and analyzing the execution performance with collaborative processes presented in the related scientific literature.
+However, their integration with process management systems is limited or non-existent.
+Against this background, we propose an approach and an architecture for blockchain-based process enactment that preserves the confidentiality of exchanged information while keeping public enforcement and transparency of process execution.
+We name our approach CONFidentiality EnforcemenT TransparencY (CONFETTY).
+In particular, we rely on smart contracts to encode and execute business process logic while logging the interactions between parties, and we resort to Multi-Authority Attribute-Based Encryption (MA-ABE) [4] to control the access of different parties to the activities’ data payloads and information artifacts, thus safeguarding confidentiality.
+We demonstrate the security of our approach against a threat model for the proposed architecture.
+We implemented our artifact and evaluated the feasibility of our solution by showing its application in the context of a healthcare scenario taken from the literature.
+Also, we analyzed the execution performance with collaborative processes presented in the related scientific literature.
 The remainder of the paper is organized as follows.
@@
 6 concludes the work and draws future research directions.
-2 Example and Problem Illustration
+2 Example, Problem Illustration and Requirements
 To illustrate the problem we aim to tackle, we introduce a running example in the healthcare domain inspired by [7].
-The model depicted in Fig.
-1 depicts the management process for X-ray diagnostics in the form of a BPMN choreography diagram [11].
+Figure 1 depicts the management process for X-ray diagnostics in the form of a BPMN choreography diagram.
 The patient first presents the medical prescription, asking for an appointment.
-The radiology department checks the availability of the ward and, if a date is available, the appointment is confirmed.
-Otherwise, there is a new scheduling tentative.
+The radiology department checks the availability of the ward, and if a date is available, the appointment is confirmed.
+Otherwise, there is a new tentative.
 After obtaining the registration details, the temperature and vaccine certification of the patient are verified.
-If the temperature is under the threshold and the certification is valid, a radiology clerk checks the appointment identifier of the patient and returns the confirmation.
-If the appointment is confirmed, the X-ray exam is done and the results are provided to the ward with a final report.
-Finally, the patient receives the results from the ward.
-As the example shows, ensuring accurate tracking of medical procedures is critical to improving process efficiency, trust in the public towards the healthcare system, and enabling auditing in case of misconduct or compliance with regulations.
-Data-based evidence is crucial to this end.
-The immutability, accessibility and persistence guarantees offered by a public-blockchain-based PAIS are fit for purpose from this standpoint.
-However, sensitive patient data must also be kept shielded from unauthorized inspections: publicly disclosing the result of the diagnostic analysis or information about vaccine certification would severely violate the privacy sphere of the patient, for example.
-Balancing a secure, publicly verifiable and decentralized infrastructure with the need for data confidentiality remains therefore an open challenge, which motivates our research.
+A radiology clerk checks whether the appointment can be confirmed.
+If so, the X-ray exam is done, and the results are provided to the ward with a final report.
+In the end, an insurance agency receives the patient’s results.
+The insurance agency can access the patient’s medical report to follow up with compensatory actions in the subscribed contract (e.g., the reimbursement of medical bills).
+Beyond the boundaries of the choreography diagram in Fig.
+1, we assume the presence of an inspector of the local Ministry of Health.
+They are not among the process participants, yet they should have full access to the information exchanged even after the conclusion of the instance for auditing purposes.
+Table 1 lists the requirements derived from the above use case.
+To begin with, ensuring public execution and tracking of healthcare procedures is critical to improving process efficiency and trust in the public towards the healthcare system.
+We concretize this aspect in Table 1.
+Enabling auditing of those procedures is crucial since it permits investigations on misconduct or lack of compliance with regulations (Table 1).
+This is allowed by the data-based evidence, which should be ensured by information storage in permanent, tamper-proof and non-repudiable information artifacts (Table 1).
+The guarantees offered by a public blockchain-based PAIS are fit for purpose from this standpoint.
+However, sensitive patient data must be shielded from unauthorized inspections (Table 1).
+For instance, although the insurance agency is involved in the process, it should not have full access to all exchanged data (e.g., the vaccine certification).
+Publicly disclosing the result of the diagnostic analysis or information about vaccine certification would severely violate the privacy sphere of the patient.
+To prevent this, access should be restricted only to designated and authorized entities through a fine-grained access control mechanism (Table 1).
+Balancing a secure, publicly verifiable and decentralized infrastructure with the need for data confidentiality remains an open challenge, which motivates our research.
+| Requirement | Approach | |
+| R1 | Public enforcement and transparency of process execution should be guaranteed | The control-flow of the process is managed by smart contracts deployed on public blockchain |
+| R2 | The process should be independently auditable with low overhead while guaranteeing authenticity | On-chain information retains publicly available process execution tracking information and hashes with locators for associated encrypted data |
+| R3 | Information artifacts should be written in a permanent, tamper-proof and non-repudiable way | Messages are stored in a tamper-proof distributed off-chain file storage, and the resource locators are stored on-chain to keep track of information |
+| R4 | Access to information artifacts should be controlled while ensuring their overall integrity and availability | MA-ABE encrypts information artifacts stored in a distributed environment.
+Decryption is possible only if the requester holds the necessary attributes.
+|
+| R5 | User-defined policies should control access levels for authenticated users with a fine-granular scheme | MA-ABE policies are associated with messages within individual activities |
 3 Background and State of the Art
-In this section, we first describe the main concepts on which our approach relies, namely blockchain technology and Multi-Authority Attribute-Based Encryption (MA-ABE), and then illustrate how those pillars constitute the foundations of existing approaches in the literature that are of relevance to our investigation.
+In this section, we first describe the core concepts underlying our approach, namely blockchain technology and Attribute-Based Encryption (ABE), and then illustrate how these pillars serve as the foundation for existing approaches in the literature relevant to our investigation.
 3.1 Background
 Blockchain technology.
-A ledger is an append-only singly-linked list of transactions, namely the transfer of assets among accounts.
-Every account is linked to a pair of keys: a public one for recognition and addressing, and a public one for the unlocking and expenditure of the account’s balance.
-A blockchain collates consecutive segments of the ledger in its body, while the header includes, among others, the number of the current block and the hash of the current and the previous block.
-The hash-link to the previous block preserves the total order of transactions in the ledger [12].
-This hash cryptographically ensures that previous blocks, and the data therein, are not changed since this would corrupt the hash-links.
-Transactions can thus be included in an append-only manner, and it is not possible to delete old interactions.
-New blocks are added to the chain at regular intervals of time by selected computational nodes within the blockchain infrastructural network.
-Those nodes are needed to compute and finalize the blocks’ hash codes, among other things.
-New cryptocurrency is minted and distributed to the nodes’ accounts to pay back their computational effort.
-The entire blockchain data structure is stored and replicated among all nodes in the network, who have full visibility over the saved ledger.
-The use of consensus protocols and cryptographic primitives permits the convergence and public verification of the transactions included in every block of each replica, thereby guaranteeing decentralization and trustworthiness without the need for an independent third-party authority certification or governance.
+A ledger is an append-only singly-linked list of transactions, representing asset transfers among accounts.
+A blockchain collates ledger segments into blocks, with headers storing block numbers and hashes linking to previous blocks, ensuring transaction order and data integrity.
+Selected nodes append new blocks and receive cryptocurrency as rewards.
+The blockchain is replicated across all nodes, ensuring transparency.
+Consensus protocols and cryptographic primitives enable public verification and decentralization without third-party oversight.
 Several platforms implement the blockchain protocol (e.g., Ethereum [31]).
-In the last few years, blockchain technologies have been adopted in many different contexts, especially after the introduction of smart contracts, namely programs deployed onto the blockchain and executed within a dedicated virtual machine (e.g., the Ethereum Virtual Machine, EVM).
-Their functions are invoked through transactions.
-Core features of smart contracts are the immutability of their code and the storage of their invocations on the blockchain, making such information available for auditing purposes [15].
-A smart contract has an account in which it can store and update data alongside its cryptocurrency balance.
-Smart contract executions incur a cost on the invoker.
+Some blockchains allow for the deployment of smart contracts [13] which are programs deployed onto the blockchain and executed within a dedicated virtual machine (e.g., the Ethereum Virtual Machine, EVM).
+Smart contracts functions are invoked through transactions and their core features are code immutability and invocation storage on the blockchain, making such information available for auditing purposes [10].
+Smart contracts execution incurs a cost on the invoker.
 The cost is measured in gas units and depends on the complexity of the called function and the size of the exchanged data.
-To save on the latter, distributed tamper-proof storage systems are often employed to store the actual data in a non-modifiable file, while a permalink to that file is saved in the smart contract state to the datastore’s locator.
+To save on the latter, distributed tamper-proof storage systems are often employed to store the actual data in a non-modifiable file, while a permalink to that file is saved in the smart contract state.
 The InterPlanetary File System (IPFS) [1] is a typical example of such a system.
@@
 The string is based on the hash of the file content itself so that if the file is altered, the permalink does not match.
-The code execution is validated in each node of the blockchain, providing verifiable proof of the obtained results for the final users [20].
 As we will see next, these features fostered the development of blockchain-based PAISs using smart contracts as the backbone for business (process) rule enforcing (in our example, the BPMN choreography is translated into smart contract code, and its steps correspond to invocations to its functions) [7, 9].
 Attribute-based Encryption.
-Attribute-Based Encryption (ABE) is a type of public key encryption where the ciphertext, namely an encrypted plaintext, and the corresponding decryption key are linked through custom boolean attributes [34].
-Ciphertext-Policy Attribute-Based Encryption (CP-ABE) [5], a variant of ABE, prescribes that every user is endowed with a set of attributes encoded in their key.
-For example, user 0xB0…1AA1 can be endowed with attributes Patient to specify their role and attribute PID476948 to indicate they are involved in process instance .
+ABE is a public key encryption where the ciphertext, an encrypted plaintext, and the corresponding decryption key are linked through custom boolean attributes [25].
+In Ciphertext-Policy Attribute-Based Encryption (CP-ABE) [2], a variant of ABE, every user bears a set of attributes encoded in their key.
+For example, the owner of the blockchain account 0xB0…1AA1 can be endowed with attributes Patient and PID476948 specifying role and process instance , respectively.
 Data owners write policies to be attached to the CP-ABE-encrypted document (the ciphertext).
@@
 Policies are propositional logic formulae employing ABE attributes as literals.
-The policy is evaluated on the attributes associated with a potential reader: Only if the formula is satisfied, the user’s key can decrypt the data.
+The policy is evaluated on attributes associated with a potential reader: Only if the formula is satisfied can the user’s key decrypt the data.
 For instance, the key of the user 0xB0…1AA1 can decrypt data locked with a policy like PID476948 and PATIENT, but cannot unlock the contents of a document associated with a policy like PID476949 and RADIOLOGY.
-Attributes are typically assigned by one authority:
+Attributes are typically assigned by one authority: Multi-Authority Attribute-Based Encryption (MA-ABE) [4] is a variant of ABE that removes this single point of failure using a multi-authority method.
+In MA-ABE, every authority creates a part of the decryption key.
+Once collected the user merges all parts to obtain the final key.
+In this paper, we employ the Ciphertext-Policy variant of MA-ABE [17].
+It integrates MA-ABE with the aforementioned policy-based approach.
 3.2 State of the Art
-Over the last few years, several solutions that automate collaborative processes using blockchain technology have been designed [13, 8].
-Several works in this field have demonstrated the effectiveness of blockchain-based solutions in improving trust among participants in multi-party collaborations [40] even in adversarial settings [26], enhance verifiability of workflows with model-driven approaches [25, 38], allow for monitoring [15], mining [21], security [23], and auditing [10, 7].
+Over the last few years, several solutions that automate collaborative processes using blockchain technology have been designed [27, 9, 18, 29].
+Such approaches demonstrated the effectiveness of blockchain-based solutions in improving trust among participants in multi-party collaborations and allowing for monitoring and auditing [10, 8].
 These studies enhance the integration of blockchain technology with process management, unlocking security and traceability benefits.
 However, they primarily focus on the control-flow perspective and lack mechanisms for secure access control to the stored data on public platforms.
-A relevant area of research involves allowing for data privacy and integrity within blockchain systems.
+A relevant area of research for our work pertains to data privacy and integrity within blockchain systems.
 Several studies examine the application of encryption techniques to achieve this goal.
 Hawk [14] employs user-defined private smart contracts to automate the implementation of cryptographic protocols.
-Abid et al.
-[1] leverage Zero-Knowledge Proof (ZKP) and Fully Homomorphic Encryption (FHE) to build a solution that supports privacy-preserving traceability and the analysis of encrypted data.
-They store on-chain only the hash of the data used to verify the correctness, while all the computations and the data are stored off-chain.
-Benhamouda et al.
-[4] use a public blockchain storage layer for confidential data.
-Peng et al.
-[32] propose a consortium blockchain as a decentralized storage system for file sharing among organizations.
-Wu et al.
-[42] present an attribute-based access control scheme that preserves the privacy of both attributes and policies.
 Pham et al.
@@
 [3] present BDABE, an access control mechanism that applies a distributed ABE scheme within a consensus-driven infrastructure for real-world applications.
-Feng et al.
-[16] combine attribute-based encryption ABE and identity-based encryption (IBE) to achieve efficient data sharing and verification of data correctness.
-They employ blockchain technology to ensure tamper-proof and regulated data storage.
 Yan et al.
@@
 The encrypted data is stored on IPFS, and the metadata is stored on the blockchain.
-All these studies, along with the aforementioned [29, 28, 27], allow for secure access control on blockchains but lack the integration of this technology with process management systems.
+All these studies, along with the aforementioned [21, 20, 19], cater for secure access control on blockchains but lack the integration of this technology with process management systems.
 To the best of our knowledge, this is the first work that focuses on combining blockchain-based process execution engines with encryption schemes that guarantee data confidentiality and access control over data shared on public blockchain platforms.
@@
 4 The CONFETTY Architecture
-In this section, we describe our proposed architecture, named CONFidentiality EnforcemenT TransparencY (CONFETTY).
-We first introduce its key functionalities and the core components that support them (Sect.
-4.1), then describe how these components interact and manage information to support confidential-preserving blockchain-based process execution (Sect.
+In this section, we describe the proposed CONFETTY architecture.
+We first introduce its core components and functionalities (Sect.
+4.1).
+Then, we describe how components interact and manage information to support confidential-preserving blockchain-based process execution while meeting Tables 1, 1, 1, 1 and 1 (Sect.
 4.2).
-Due to space restrictions, we do not enter the technical details of our approach but provide an overarching presentation of our solution’s rationale and key concepts.
-Finally, we conclude the section with some considerations on the benefits brought by our approach merging blockchain technologies and ABE.
+Due to space restrictions, we do not enter the technical details of our approach but provide an overarching presentation of our solution’s rationale and key concepts, motivating them with the requirements they aim to meet.
 4.1 Functional Viewpoint
-Figure 2 outlines the CONFETTY architecture,
+Figure 2
+provides an overview of the CONFETTY approach,
 which caters for the following main functionalities.
-(i) The configure functionality foresees the specification of the process by the process owner.
-It requires the registration of process information including activities to execute, data to exchange, control-flow routing, the expected participants’ roles (e.g., the information enclosed in the choreography diagram in Fig.
-1) and how they can operate with confidential data (via ABE policies) to control information access.
-(ii) The instantiate functionality lets certifiers authenticate the participants, define their role in the process (e.g., that user 0xB0…1AA1 is a patient), and also kickstart the process run on-chain, by setting its public state.
-(iii) The transact functionality allows process participants to interact following the process to exchange data at run time.
+(i) The configure functionality helps the process owners specify the process to run on-chain.
+It requires the registration of process information, including activities to execute, data to exchange, control-flow routing, the expected participants’ roles (like the information enclosed in the choreography diagram in Fig.
+1) and how they can operate with confidential data (via MA-ABE policies) to control information access.
+(ii) The instantiate functionality lets certifiers authenticate the participants, define their role in the process (e.g., that user 0xB0…1AA1 is a patient), and kickstart a new process instance on-chain by setting its public state.
+(iii) The transact functionality allows process participants to interact, following the process, to exchange data at run time.
 This functionality manages both public and confidential data, updating and enforcing the process state according to the logic in the blockchain.
 (iv) The inspect functionality allows participants and auditors to access public and confidential data.
-The functionalities above are realized by four main components of CONFETTY: (i) The Process Manager handles all the operations involving the participants and related to the process.
+The functionalities above are realized by the CONFETTY main components depicted in Fig.
+3.
+CONFETTY relies on two external components used as architectural buttresses: (i) A Programmable Blockchain Platform (like Ethereum), (ii) and a Tamper-proof Distributed File Storage (like IPFS).
+The former maintains the public state, allowing for the creation of logic and constraints while notarizing the executed operations.
+It comprises two core elements: The Distributed Ledger, storing the transactions, and the Virtual Machine, hosting and running smart contracts.
+The latter (which we will henceforth name as Distributed FS for brevity) stores business data and thus improves the performance and costs of a blockchain as in common practice.
+The data is saved on the external storage (upon encryption, if it must remain secret), while a resource locator with the hash of that data is kept on-chain for notarization and future retrieval.
+Notice that we take inspiration from known patterns adopting Distributed FSs for large data storage [32].
+Based on this known pattern, we devise new techniques to preserve secrecy and control access to that process data at runtime.
+The primary components of CONFETTY are the following: (i) The Process Interface handles all the operations involving the participants and related to the process.
 This is done by acting as an interface for the blockchain, abstracting from the technical implementation and exposing functionalities to the users.
-In particular, the Process Manager instantiates the process on the blockchain on given process specification data.
+In particular, the Process Interface instantiates the process on the blockchain on given process specification data.
 To support the instantiate functionality, it provides capabilities for role assignment to participants.
 (ii) The Process Contract is a smart contract that manages the process instances and their execution.
-It elaborates and manages all process specifications and their instances’ public states, acting as a software realization both of factory and proxy patterns [17] akin to [10].
+It elaborates and manages all process specifications and their instances’ public states, acting as a software realization both of factory and proxy patterns [32].
 It supports the instantiate functionality by enforcing the control flow, the data flow, the assignment of tasks to participants, and the routing of decision points.
 In this way, the Process Contract updates the public state of the process and generates transparent records that are accessible and verifiable by interested auditors.
-(iii) The Confidentiality Manager is in charge of handling the participants’ authorizations and the operations over confidential data.
-It operates as a facade towards the blockchain to store access grants, distributes decryption keys to the users solely based on their roles and involvement in the process, and handles the storage and retrieval of encrypted data.
+(iii) The Confidentiality Interface is in charge of handling the participants’ authorizations and the operations over confidential data.
+It operates as a façade towards the blockchain to store access grants, distributes decryption keys to the users solely based on their roles and involvement in the process, and handles the storage and retrieval of encrypted data.
 (iv) The Confidentiality Contract is a smart contract handling the notarization of authorizations for confidential data, as well as for their writing and access.
-In addition to the above inner components, the architecture relies on two external ones: a Programmable Blockchain Platform (like Ethereum) and a Tamper-proof Distributed File Storage (like IPFS).
-The former maintains the public state, allowing for the creation of logic and constraints while notarizing the executed operations.
-It comprises two core elements: The Distributed Ledger, storing the transactions, and the Virtual Machine, hosting and running smart contracts.
-The latter (which we will henceforth name as Distributed FS for brevity), is used to store business data and thus improve the performance and costs of a blockchain.
-The data is saved on the external storage (upon encryption, if it must remain secret), while a resource locator with the hash of that data is kept on-chain for notarization and future retrieval.
-We conclude this subsection with a few considerations about the actors and trust model.
-We assume that the aforementioned users (process owners, certifiers, process participants, and auditors) possess a blockchain account with a key pair to be able to publicly identify themselves (as with the aforementioned user 0xB0…1AA1) and sign blockchain transactions built through the CONFETTY modules.
+We conclude this subsection with a few considerations about the users and trust model.
+We assume that the users we mentioned while discussing the functionalities (i.e., process owners, certifiers, process participants, and auditors) possess a blockchain account with a key pair to be able to publicly identify themselves (as with the aforementioned user 0xB0…1AA1) and sign blockchain transactions built through the CONFETTY modules.
 We also remark that we consider the CONFETTY platform as a trusted party, while users are assumed to be honest but curious.
 Finally, notice that the involvement of a pre-appointed third-party certifier to authenticate process participants is necessary as a user alone can confirm their identity (e.g., via the aforementioned public-private key scheme) but cannot attest alone to the truthfulness of assertions regarding themselves.
+The selection and appointment of certifiers transcends the goals of this paper, but in our motivating scenario, e.g., certifiers can be trusted third parties like the Public Ministry of Health (for hospital personnel), the Public Registry (for citizens), and the Insurance Registrar (for agencies), or subsidiaries thereof.
+In our architecture, certifiers operate like blockchain oracles [23].
 4.2 Information Viewpoint
 After providing an overview of the CONFETTY architecture, we describe how information is handled in realizing the aforementioned functionalities by the CONFETTY components.
-Figure 3 depicts the configure functionality.
-The process owner sends the process specification files (like the diagram in Fig.
-1) to the Process Manager (1.0).
-The Process Manager, in turn, (1.1) processes the input data, storing the behavioral and business constraints it reports.
-Also, (1.2) it builds a policy template, associating roles of the process with the data they can access via encryption directives.
-Once created, the policy template is returned to the process owner.
-In our running example, an extract of a policy template is:
-{medicalPrescription:($PID and (PATIENT or RADIOLOGY)),
-registration:
-($PID and (RADIOLOGY or PATIENT)), report:($PID and (RADIOLOGY or WARD))}.
-Here medicalPrescription, registration, and report represent messages, while PATIENT, RADIOLOGY and WARD are the roles of users who can access them.
-$PID is a placeholder for the process instance identifier (a piece of information that can be known only at runtime).
-We give the process owner the option to revise the templates.
-In this way, adding custom roles that are not expected active parties in the process but need access to data.
-This is the case, for instance, of external auditors who need to certify data and its compliance with norms, laws and regulations.
-Table 1 shows an extract from the final access policies of our running example.
-In this case, each message is attached to a policy with an ID representing the process instance and the authorized role.
-Considering the medicalPrescription message, the policy specifies that to access it, a user has to be attested to have the role of a MINISTRY-INSPECTOR, or again to be a PATIENT or RADIOLOGY role within the running process instance.
 | Message | Access policy |
@@
 | Report | MINISTRY-INSPECTOR or ($PID and (RADIOLOGY or WARD)) |
-Once policies are confirmed, (1.3) they are sent to the Confidentiality Manager, which (1.4) stores them in the Distributed FS.
-Their permalink (1.5) is passed back to the Confidentiality Manager which saves it.
+The first functionality is configure.
+Here the process owner sends the process specification files (like the diagram in Fig.
+1) to the Process Interface.
+The Process Interface, in turn, processes the input data, storing the behavioral and business constraints it reports.
+Also, it builds a set of parametric access policies, associating roles of the process with the data they can access via encryption directives (meeting Table 1).
+Once created, the policies are returned to the process owner.
+The one associated with the medical prescription (sent by the patient to the radiology clerk to take an appointment in Fig.
+1), e.g., reads “$PID and (PATIENT or RADIOLOGY)”.
+Here, PATIENT and RADIOLOGY are the attributes representing the actors’ roles in the process.
+Not all patients and radiology clerks should access any medical prescription.
+Hence $PID, an attribute placeholder for the process instance identifier (a piece of information that can be known only at runtime), restricting access to the sole actors involved in that process instance.
+We give the process owner the option to revise the policies by adding custom roles that are not expected active parties in the process but need access to data (as per Table 1).
+This is the case, for instance, of external auditors who need to certify data and its compliance with norms, laws and regulations.
+Considering the medical prescription, the revised policy specifies that a user has to be attested to have the role of a MINISTRY-INSPECTOR, or again to be a PATIENT or RADIOLOGY within the running process instance: “MINISTRY-INSPECTOR or ($PID and (PATIENT or RADIOLOGY))”.
+Table 2 shows an extract from the final access policies of our running example.
+Once policies are confirmed, they are sent to the Confidentiality Interface, which stores them in the Distributed FS.
+Their permalink is passed back to the Confidentiality Interface to save it.
 Figure 4 depicts the instantiate functionality.
-Differently from the configure functionality, needed once per process, this one is designed for every new instantiation thereof.
-To initialize a new process instance (e.g., number 476948), participants send their data to the Process Manager (2.0) to authenticate themselves as participants in the process with their role (e.g., that user 0xc5…Fd43 operates as a WARD clerk in a new process instance PID476948).
-(2.1) The authentication is attested to by certifiers confirming the user’s self-declared roles.
-When all process participants are identified and associated with a role, (2.2) the Process Manager merges the received information with the process data acquired in step 1.1.
-Thereupon, (2.3) a selected participant kickstarts a new process instance through the Process Manager associated with the aforementioned information on the Process Contract, thereby initializing the instance’s public state.
-Notice that also this crucial step is notarized on chain (2.4).
+Differently from the configure, needed once for each process, this one is designed for every new instantiation thereof.
+To initialize a new process instance (e.g., number 476948), participants send their data to the Process Interface (1) to authenticate themselves as participants in the process with their role (e.g., that user 0xc5…Fd43 operates as a WARD clerk in a new process instance PID476948).
+(2) The authentication is attested to by certifiers confirming the user’s self-declared roles.
+When all process participants are identified and associated with a role, (3) the Process Interface merges the received information with the process data acquired in the configuration.
+Thereupon, (4) a selected participant kickstarts a new process instance through the Process Interface associated with the aforementioned information on the Process Contract, thereby initializing the instance’s public state.
+This step is done by sending a transaction to the Process Contract with the public state of the process, including conditions driving the control flow, such as exclusive gateway choices (as per Table 1).
+Notice that this crucial step is notarized on chain (5) (meeting Table 1).
 The selection of the participant, or participants, on behalf of whom the transaction is signed, transcends the scope of this paper and depends on the decision process and guidelines of the consortium.
-After initializing the process state, (2.5) the Process Manager triggers the Confidentiality Manager, which (2.6) stores the policy locator previously saved on the Confidentiality Contract.
-Finally, (2.7) the Confidentiality Manager notarizes the operation.
-The transact functionality reported in Fig.
-5 represents the exchange of information during the execution of an activity in the process instance.
-This passage updates the public state update in the blockchain according to previous enforcement (e.g., activity to execute, expected participant, etc.) and manages accessibility in case of public and confidential cases.
-The first case refers to public data and is displayed in Fig. 5(a).
-Initially, (3.0a) a user sends the execution data to the Process Manager in the context of an activity. For example, the WARD clerk identified by 0xc5…Fd43 replies to the
+After initializing the process state, (6) the Process Interface triggers the Confidentiality Interface, which (7) stores the policy locator previously saved on the Confidentiality Contract.
+Finally, (8) the Confidentiality Interface notarizes the operation.
+Figure 5 displays the transact functionality, which represents the exchange of information during the execution of an activity in the process instance.
+It encompasses a public and a confidential case, depending on the process participants and their willingness to classify information as disclosable or not.
+In both cases, this functionality updates the public state in the blockchain according to previous enforcement as per Table 1 (e.g., activity to execute, expected participant) and manages accessibility (meeting Table 1).
+Let us begin with the exchange of public data (and the update of the public state). The main steps are highlighted in blue in Fig. 5 and salient on-chain operations are reported in Algorithm 1.
+Initially, a user sends the execution data to the Process Interface in the context of an activity (Table 1).
+For example, the WARD clerk identified by 0xc5…Fd43 replies to the
 “Check availability” activity request by the RADIOLOGY user within process instance 476948.
@@
 In our scenario, the exclusive choice gateway is based on the value of the accepted variable.
-Then, (3.1a) the Process Manager invokes the Process Contract to request a state update, sending the execution data.
-In turn, (3.2a) the Process Contract enforces the business and control-flow logic through the run of its routine (based on previously agreed upon and immutable code, see step 1.1), which checks if the executed activity is enabled and if the message sender (in this case, user 0xc5…Fd43) is the expected one (a WARD), finally updating the state if the verification outcome is positive.
-Also, (3.3a) this step advances the state of the process, updating the expected next activity (here, “Confirm appointment”) and participant (RADIOLOGY).
-This piece of information is notarized into the Distributed Ledger.
-The confidential case is depicted in Fig.
-5(b).
-The Process Manager, (3.0b) once it has received the data, (3.1b) forwards it to the Confidentiality Manager.
-The latter (3.2b) retrieves the policy from the Distributed FS, and (3.3b) uses it to encrypt the data via Attribute-Based Encryption (ABE).
+Then, the Process Interface invokes the Process Contract to request a state update, sending the execution data (Line 1).
+In turn, the Process Contract enforces the business and control-flow logic through the run of its routine (based on previously agreed upon and immutable code, see configure functionality), which checks (Line 1) if the executed activity is enabled and if the message sender (in this case, user 0xc5…Fd43) is the expected one (a WARD), finally updating the state (Line 1) if the verification outcome is positive.
+This step advances also the state of the process (Line 1), updating the expected next element (here, “Confirm appointment”) and participant (RADIOLOGY).
+In case the element is a gateway, its logic is automatically executed to advance the control flow of the process (Line 1).
+In the end, this piece of information is notarized into the Distributed Ledger (meeting Table 1).
+For the confidential case, we refer again to Fig.
+5 (see the marks in red).
+Overall, (1) once the Process Interface has received the data, (2) it forwards it to the Confidentiality Interface.
+The latter (3) retrieves the policy from the Distributed FS, and (4) uses it to encrypt the data via MA-ABE (as per Table 1).
 For example, considering Fig.
 1 and Table 2, the medical prescription of the patient is encrypted with the following policy: {MINISTRY-INSPECTOR or (PID476948 and (PATIENT or RADIOLOGY))}.
-In particular, the Confidentiality Manager retrieves the specific authorization for the data to be stored and embeds it in the encryption mechanism, ensuring that only authorized participants can access it.
-Once encrypted, (3.4b) the Confidentiality Manager stores data in the Distributed FS, (3.5b) its locator in the Confidentiality Contract, and (3.6b) notarizes the operation in the Distributed Ledger.
+In particular, the Confidentiality Interface retrieves the specific authorization for the data to be stored and embeds it in the encryption mechanism, ensuring that only authorized participants can access it.
+This step is needed only in confidential cases, as in public ones, information is directly stored on the blockchain and intentionally kept publicly readable.
+Once encrypted, (5) the Confidentiality Interface stores data in the Distributed FS, (6) its locator in the Confidentiality Contract, and (7) notarizes the operation in the Distributed Ledger (meeting Table 1 and Table 1).
 In this case, the resource locator that contains the encrypted medicalPrescription (e.g., LXBw…2DAx) is Qmo9…0weI.
-The execution of this routine by the Confidentiality Contract includes a call to the Process Contract (3.7b) to request a state update (enclosing the execution data).
+The execution of this routine by the Confidentiality Contract includes a call to the Process Contract (8) to request a state update enclosing the execution data.
 Following the example, this data is the activity being executed (“Check availability”), the user that performs it (e.g.,, PATIENT, along with their corresponding address 0x2e…6dd9), and the process instance (e.g.,476948).
-Then, (3.8b) the Process Contract checks the current public state to update it. The operations involving the smart contracts (3.9b) are notarized in the Distributed Ledger.
+Then, (9) the Process Contract checks the current public state to update it. The operations involving the smart contracts (10) are notarized in the Distributed Ledger.
 The inspect functionality pertains to the access to message data.
-The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above).
-To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester.
+The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above) (Table 1).
+To read public data, a process participant or an auditor may send a request to the Process Interface, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester.
 This is the case, e.g., for the accepted field of the “Check availability” task.
-Reading confidential data requires more passages, illustrated in Fig.
-6, and requires a preliminary step.
-All users should request at least once (and at every attribute update) a decryption key (Fig.
-6(a)), to be used for every message they are going to read (Fig.
-6(b)).
-For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a).
-The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger.
-Then, (4.3a) the Confidentiality Manager generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via Attribute-Based Encryption (ABE), and (4.4a) sends it back.
-The decryption of confidential data is depicted in Fig.
-6(b).
-Once the process participant has received the a-b key, they can use it to decrypt the confidential data.
-To do so, (4.0b) the user makes a request to the Confidentiality Manager.
-Then, (4.1b) the Confidentiality Manager retrieves the encrypted data from the Distributed FS and (4.2b) forwards it to the user.
-At this point, the latter (4.3b) uses the a-b key to decrypt the confidential data via ABE.
-Only if the user’s attributes satisfy the policy previously used to encrypt the data, can the user’s a-b key decrypt the confidential data (and the user inspect it in-clear).
+Reading confidential data requires more passages and a preliminary step.
+All users should request at least once (and at every attribute update) a decryption key, to be used for every message they are going to read.
+For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Interface.
+The latter forwards the request to the Confidentiality Contract, which is thus notarized on the Distributed Ledger.
+Then, the Confidentiality Interface generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via MA-ABE, and sends it back.
+The aforementioned step enables the decryption of confidential data.
+Once the process participants have received the a-b key, they can use it to decrypt the confidential data.
+To do so, a user makes a request to the Confidentiality Interface.
+Then, the Confidentiality Interface retrieves the encrypted data from the Distributed FS and forwards it to the user.
+At this point, the latter uses the a-b key to decrypt the confidential data via MA-ABE.
+Only if the user’s attributes satisfy the policy previously used to encrypt the data can the user’s a-b key decrypt the confidential data (and the user inspects it in clear).
 A note on decryption keys.
-We remark that an a-b key stays with the user and lasts for as long as their attributes are not updated (in case, e.g., a user is involved in a new process instance, say 476949, a new attribute gets associated to them: $PID476949).
+We remark that an a-b key stays with the user and lasts for as long as their attributes are not updated (if a user is involved in a new process instance, say 476949, a new attribute gets associated with them: PID476949).
 The same key will be used to access all messages.
 For example, consider user 0xB0…1AA1 (who holds the attributes PID476948 and PATIENT).
-As per the policies in Table 1, their a-b key can decrypt both the medical prescription and the registration alike, but not the report.
-The a-b key of user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
+Per the policies in Table 2, their a-b key can decrypt the medical prescription and the registration, but not the report.
+The a-b key of the user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
 Notice that a HEALTH-INSPECTOR, instead, can use their unique a-b key to access all the aforementioned documents regardless of the process instance they pertain to.
-The key, in other words, is associated to the user and it is the sole instrument to access a document.
-Therefore, the Confidentiality Manager does not need to generate a different decryption key for every document and give the same copy to multiple users, nor create multiple copies of the same document for different decryption keys (one per user), thus saving on unnecessary data replications and key distributions.
-Also, as the key stays with the users, and the documents are stored and notarized on tamper-proof external infrastructures (the Distributed FS and Blockchain Platform, respectively) the information remains available even should the CONFETTY system stop functioning.
-These aspects have an impact on security considerations, as we will discuss next.
+The key, in other words, is associated with the user, and it is the sole instrument to access a document.
+Therefore, the Confidentiality Interface does not need to generate a different decryption key for every document and give the same copy to multiple users, nor create numerous copies of the same document for different decryption keys (one per user), thus saving unnecessary data replications and key distributions, and ensuring integrity (Table 1).
+Also, as the key stays with the users, and the documents are stored and notarized on tamper-proof external infrastructures (the Distributed FS and Blockchain Platform, respectively), the information remains available even should the CONFETTY system stop functioning (Table 1).
+These aspects have an impact on security considerations, as we discuss next.
 5 Evaluation
-To assess the feasibility of the CONFETTY architecture we implemented a supporting prototype and evaluated it.
-We first describe the implemented architecture analyzing its thread model, then we evaluate CONFETTY in operation.
-5.1 Implementation
-Figure 7 illustrates the technologies utilized in the implementation of CONFETTY.
-Our prototype is deployed on three main tiers.
-An application server hosts the Confidentiality Manager and the Process Manager, developed in Python 3.6.9 and Java SDK-13, respectively.
-Users communicate with it via TLS communication channels [37].
-We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache.
-Accessed: December 13, 2024.
-as the Distributed Tamper-proof File Storage.
-We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js.1 The source code of CONFETTY, alongside the experimental data we use, is publicly available at github.com/apwbs/CONFETTY.
+To assess the feasibility of the CONFETTY architecture, we perform a two-step evaluation.
+First, we analyze the CONFETTY threat model, evaluating it against the proposed architecture.
+Then, we describe the implemented prototype and analyze its performance.
+The source code of CONFETTY, alongside the experimental data we use and performance analysis results, are available at https://doi.org/10.5281/zenodo.15482587.
 5.1 Threat Model
-In this section, we identify potential security threats that could compromise data confidentiality within our solution, and outline how the CONFETTY architecture addresses them.
+In this section, we identify potential security threats that could compromise data availability (Table 1) and confidentiality (Table 1) within our solution, and outline how the CONFETTY architecture addresses them.
 Our threat analysis is based on the STRIDE framework [26], a theoretical model that categorizes threats into six distinct groups: spoofing (i.e., impersonation of a legitimate entity), tampering (i.e., unauthorized modification of data to compromise its integrity), repudiation (i.e., the denial of having performed a particular action), information disclosure (i.e., the unauthorized exposure of sensitive data), denial of service (DoS, i.e., the disruption or degradation of system availability), and elevation of privileges (i.e., the unauthorized acquisition of higher-level access rights).
-As said in Sect.
-4.1, we assume users and components to be honest but curious and trusted, respectively.
-The possible threats thus reside in the communication between them, subject to potential attacks.
+We assume users and components to be honest but curious and trusted, respectively.
+The possible threats thus reside in the communication between them, subject to potential attacks (see Sect.
+4.1).
 Figure 6 displays a Data Flow Diagram (DFD), in which the orange dotted lines indicate the trust boundaries crossing data exchange points between different perimeters, represented by black arrows connecting different elements.
@@
 The CONFETTY Boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditor).
-In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented.
+In this case, a spoofing attack could result in a user impersonating another.
+Still, this attack is prevented since we assume that each user is linked to the Blockchain Platform credentials.
 As for tampering, trust is based on the fairness of the user who sends the data.
@@
 An attacker could sniff the data a user sends to CONFETTY.
-However, for those communications we resort to TLS channels, specifically designed to prevent information leakage [37].
-DoS attacks are in principle feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform.
-Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1.0 in Sect.
-4.2) which malicious nodes could exploit with multiple calls and bulky payloads.
-However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss.
-Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE.
-Elevating a user’s privileges is not possible since attributes and permissions are linked to a blockchain account.
+However, we resort to TLS channels for those communications, specifically designed to prevent information leakage [28].
+DoS attacks are, in principle, feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform.
+Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1 in Fig.
+5) which malicious nodes could exploit with multiple calls and bulky payloads.
+However, DoS attacks would lead to a disruption or degradation of the system’s availability.
+Yet, all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and the Distributed Tamper-proof File Storage to reactivate or replicate the system without any data loss.
+Information disclosure is prevented because all sensitive information is encrypted via MA-ABE.
+Elevating a user’s privileges is impossible since attributes and permissions are linked to a blockchain account.
 The Blockchain Platform is the trusted component in the Blockchain Boundary (BB).
-In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the real one.
-This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their own private keys.
+In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the original software object.
+This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their private keys.
 Information disclosure can be achieved only on public data or resource locators of sensitive data.
-However, the latter is stored upon ABE-encryption, making its content unreadable by malicious users.
+However, the latter is stored upon MA-ABE-encryption, making its content unreadable by malicious users.
 Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols.
-As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack, and elevation of privileges, are counteracted by design, too.
+As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack and elevation of privileges are counteracted by design, too.
 Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain.
-Since we resort to ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.
+Since we resort to MA-ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.
+5.2 Implementation
+CONFETTY is deployed on three main tiers.
+An application server hosts the Confidentiality Interface and the Process Interface, developed in Python 3.6.9 and Java SDK-13, respectively.
+Users communicate with it via TLS communication channels [28].
+We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: web3js.readthedocs.io/en/v1.10.0/.
+Accessed: July 16, 2025.
+as the Distributed Tamper-proof File Storage.
+We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js1.
 5.3 Execution Analysis
 […]
-Table 2 shows the execution time and costs registered during the execution of the X-ray diagnostic process.
-For the time evaluation, we ran a separate process instance for each path in the control flow (each corresponding to a different decision in correspondence with the exclusive-choice gateways) five times (thus totalling twenty executions).
-The registered values are reported in Table 2(a).
-We employed the Sepolia testnet as the Blockchain Platform, and passed input data based on real-world files in the healthcare domain upon anonymization.
-The dataset and the Sepolia transaction handles are openly accessible at github.com/apwbs/CONFETTY.
-[…]
-| min | max | avg | |
-| Configure | |||
-| Instantiate | |||
-| Transact | |||
-| Inspect | |||
-| Functionality | Used gas units |
-| Instantiate | |
-| Transact | |
-| Inspect | |
-| Deployment | |
-| Total |
+5.3.1 Execution Performance and Expenditure.
+We gauged the time and cost of our prototypical implementation against data simulating the execution and message exchange of processes from the literature.
+Figure 7 illustrates the results, averaging over runs that took the longest path.
+We gathered data using a local RPC as a stub for the Programmable Blockchain Platform.
 […]
 […]
-The second part of our experimental evaluation illustrates the time and cost needed by CONFETTY to run three different processes from the literature, namely the aforementioned X-ray [9], retail [7] and incident management [10].
-Figure 9 illustrates the results, averaging overruns that took different branches in correspondence with exclusive choice gateways.
-We gathered data using a local RPC as a stub for the Programmable Blockchain Platform.
-The three processes exhibit similar trends from both standpoints.
-The Incident management process is the heaviest and most expensive one, yet exhibiting an increment of for the cumulative time and for the cumulative cost with respect to the lower bound of the X-ray process.
-This increment depends on the number of process participants ( in the Incident management against the of the X-ray process), thus leading to the execution of more steps.
+Let us now compare the above results with two other processes from the literature, namely retail [6] and incident management [8].
+The X-ray, incident and retail processes have similar structures in the number of messages (10, 13 and 12, respectively), gateways (5, 3 and 2) and involved participants (4, 5 and 3) and exhibit similar trends from both runtime and cost standpoints.
+The incident management process is the slowest and most expensive one with an increase in cumulative time and cost compared to the X-ray process, respectively.
+The experiments show that CONFETTY performs comparably to literature tools and scenarios.
+Indeed, ChorChain [8] required approximately million GU for the incident management process, while MultiChain [6] required around million GU for the retail process.
+Notice that such approaches only focus on public execution and enforcement without incorporating any confidential mechanisms and that they deploy a contract for each new instance.
+5.3.2 Scalability Assessment.
+Lastly, we analyzed the scalability of CONFETTY.
+To this end, we considered the dependence of time and cost taking as independent variables the dimensions of the process affecting the input size for our tool.
+In this section, we focus on the number of process participants and the choreography model size due to space restrictions as they evidence the most salient results.
+Further experiments including other variables like the number of gateways (parallel and exclusive) and the message payload size can be found in the online repository.
+The default values for the variables adhere to the example provided in Fig.
+1.
+For every test, we let one variable change the value and keep the others fixed, assigned with the respective defaults.
+Figures 8 and 9 dissect the registered timings and costs to serve the different functionalities concurring to the total by stacking the respective subtotals.
+[…]
+displays the results observed varying the process model size.
+We progressively increased it by sequentially concatenating the whole choreography diagram multiple times to derive a new one, thus scaling the size up to ten times the original one.
+Figure 9(a) shows a linear growth in the runtime of the transact and inspect functionalities, without a significant impact on the remaining one, instantiate.
+The execution time increases with the number of messages due to additional read and write operations.
+Figure 9(b) shows the cost increase, with a linear growth registered for the instantiate and transact functionalities.
+More elements increase on-chain storage needs and hence gas cost.
+Read operations remain constant as they only handle one-time key management on-chain (hence the flat cost associated with inspect).
 6 Conclusions
 In this work, we presented CONFETTY, a blockchain-based architecture for PAISs.
-CONFETTY allows for transparency and public enforcement of process execution while preserving confidentiality of sensitive data.
-A public state saved on chain maintains process data used for enforcement and auditability purposes.
-The evolution of the public state is mediated by smart contract implementing the business process control- and data-flow logic.
-We use ABE encryption to hide sensitive information and grant access to it only to authorized parties based on their roles in the process, encoded as attributes in access policies and embedded directly in their personal decryption keys.
-To evaluate the CONFETTY architecture we have analyzed its threat model and measured its time and cost performance through the execution of an X-ray examination choreography process.
-Future work includes extensive performance and scalability analyses against synthetic and real-world benchmarks.
-We also envision a field study to assess the practical adoption and usability with real-world stakeholders.
-Another interesting avenue for further investigations is the full decentralization of the architecture, to remove the existing points of trust residing in off-chain components.
-Finally, we foresee a study and design of a solution to enable automated decision support via computation held on encrypted data.
+CONFETTY enables transparency and public enforcement for process execution while preserving the confidentiality of sensitive exchanged data.
+A public state saved on-chain maintains process data for enforcement and auditability purposes.
+The evolution of the public state is mediated by smart contracts implementing the business process control- and data-flow logic.
+We use MA-ABE encryption to hide sensitive information and grant access only to authorized parties based on their roles, encoded as attributes in access policies and embedded directly in their decryption keys.
+We confirmed its security guarantees with a threat model analysis, and demonstrated the time and cost performance of our implemented prototype with processes taken from the literature.
+In future work, we envision a field study to assess the practical adoption and usability with real-world stakeholders.
+Another interesting avenue is the full decentralization of the architecture (to remove the existing points of trust residing in off-chain components) and governance (to cater for multiple process owners during the instantiation phase).
+Conducting a security analysis of our approach against stronger adversarial models is an intriguing future endeavor, too.
+Finally, we envision a study and design of a solution to enable automated decision support via computation held on encrypted data.
+6.0.1 Acknowledgements
+This work was partly funded by projects PINPOINT (B87G22000450001) under the PRIN MUR program, and Health-e-Data, funded by the EU-NGEU under the Cyber 4.0 NRRP MIMIT programme.